(stock issuances and repurchase)

Lindsey Hunter Corporation is authorized to issue 50,000 shares of $5 par value common stock. During 2017, Lindsey Hunter took part in the following selected transactions.

1) issued 5,000 shares of stock at $45 per share, less costs related to the issuance of the stock totaling $7,000

2) issued 1,000 shares of stock for land appraised at $50,000. The stock was actively traded on a national stock exchange at approximately $46 per share on the date of issuance.

3) Purchased 500 shares of treasury stock at $43 per share. The treasury shares purchased were issued in 2013 at $40 per share.

Instructions:

A) Prepare the journal entry to record item 1

B) Prepare the journal entry to record item 2

C) Prepare the journal entry to record item 3 using the cost method.
Answer #1


Solution: No. Account title Cash =5000*45-7000 Common Stock =5000*5 Paid-in Capital in Excess of Par-Common Stock Debit Credi